value of [(2014^2-2020) (2014^2+4028-3) (2015)] / [(2011) (2013) (2016) (2017)], ()

Grade:10

4 Answers

Mimi Bhattacharya
5 years ago
2014^2+4028-3
=2014^2+4028+1-4
={2014^2+(2×2014×1)+1^2}-4
=(2014+1)^2- 2^2
=(2014+1+2)(2014+1-2)
=2017×2013
 
2011×2016
=(2014-3)(2014+2)
=2014^2-(3+2)2014-(3×2)
=2014^2-2020
 
When we'll arrange the numerator and denominator, it'll be noticed that nearly all the numerator are cancelled and we're left with only 2015.
Thus, the answer is 2015.
daksh
3 years ago
Here what I found the best and easy explanation on it ever,
you must check it out
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
..
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
..
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
..Here what I found the best and easy explanation on it ever, you must check it out
..
Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out Here what I found the best and easy explaination on it ever,you must check it out
daksh
3 years ago
Here what I found the best and easy explanation on it ever,
you must check it out
add capital Y to the last of the link
htt
daksh
3 years ago
Here what I found the best and easy explanation on it ever,
you must check it out
search it on google and paste it and click on the first link
q2cLJBfRnZY